### v2

@@ 9 Conclusion
We precisely defined the concept of next-token prediction capacity, proved general upper bounds, and proved a lower bound for one-layer multi-head decoder-only transformers. Assuming that the activation function is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters are necessary to memorize a data set with $n$ unique contexts and a vocabulary of size $\omega$. Assuming that the activation function is real analytic at a point and not a polynomial there (e.g. arctan, tanh, GELU), we showed that $\Omega(n\omega)$ parameters is also sufficient. Furthermore, we provided numerical evidence that the transformer can already be trained to the entropy lower bound in the $\Theta(n\omega)$ parameter regime. Theoretically investigating optimization and generalization in the $\Theta(n\omega)$ parameter regime is left as a future direction. It is also left as a future direction to analyze the next-token prediction capacity for multi-layer transformers and for modified architectures, such as Mixture of Softmaxes, which are not back-loaded, i.e. which have a greater proportion of parameters in the embedding and self-attention sub-layers than in the FNN sub-layer. Finally, it seems worthwhile to consider memory capacity in the more constrained setting of stationary processes.

### v3

@@ 9 Conclusion
We precisely defined the concept of next-token prediction capacity, proved general upper bounds, and proved a lower bound for one-layer multi-head decoder-only transformers. Assuming that the activation function is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable (e.g. arctan, tanh, GELU), we showed that $\Omega(n\zeta)$ parameters are necessary to memorize a data set with $n$ unique contexts and a vocabulary of size $\zeta$. Assuming that the activation function is real analytic at a point and not a polynomial there (e.g. arctan, tanh, GELU), we showed that $\Omega(n\zeta)$ parameters is also sufficient. Furthermore, we provided numerical evidence that the transformer can already be trained to the entropy lower bound in the $\Theta(n\zeta)$ parameter regime. Theoretically investigating optimization and generalization in the $\Theta(n\zeta)$ parameter regime is left as a future direction. It is also left as a future direction to analyze the next-token prediction capacity for multi-layer transformers and for modified architectures, such as Mixture of Softmaxes, which are not back-loaded, i.e. which have a greater proportion of parameters in the embedding and self-attention sub-layers than in the FNN sub-layer. Finally, it seems worthwhile to consider memory capacity in the more constrained setting of stationary processes.